mathslover (mathslover):

I just forgot the way to find the rationalizing factor can any1 help me with this?

mathslover (mathslover):

I just want an example for this @Ganpat @jagatuba and @saifoo.khan

OpenStudy (anonymous):

okay i will

mathslover (mathslover):

Hmn please go ahead @Nick_Black

OpenStudy (anonymous):

http://www.twiddla.com/895293 come here

mathslover (mathslover):

You there?.

mathslover (mathslover):

@saifoo.khan can u help/?

OpenStudy (saifoo.khan):

Isn't @Nick_Black helping?

OpenStudy (anonymous):

aap hain yaha pe?? @saifoo.khan

mathslover (mathslover):

But i didn't get how to get rationalizing factor @Nick_Black ..

OpenStudy (anonymous):

u rationalize the denominator this is to change it frm sqrt to a whole no.

OpenStudy (ganpat):

@mathslover : Rationalisation is multiply and divide by denominator... so as to remove any square root sign in the denominator... This is just to simplify further calculations..

OpenStudy (ganpat):

And the denominator is denoted as factor in the equation.. SO rationalizing factor..

OpenStudy (anonymous):

ganpat is righjt here

mathslover (mathslover):

Oh k!!! thanks a lot

OpenStudy (lgbasallote):

a summary: \[\frac{a}{\sqrt b} \implies \frac{a}{\sqrt b} \times \frac{\sqrt b}{\sqrt b} \implies \frac{a\sqrt b}{b}\]

OpenStudy (anonymous):

Nice pic @lgbasallote

OpenStudy (jiteshmeghwal9):

is this nice photo dude

OpenStudy (anonymous):

This is the case when denominator is a single term: If it is of two terms then: We always change the sign of one out of them which you can think can make my denominator simple.. For example: Suppose you have: \[\frac{a}{x - \sqrt{y}}\] Change the sign of \(\sqrt{y}\) so that you can use : \((m-n)(m+n) = m^2 - n^2\) So you will multiply and divide by \(x + \sqrt{y}\) ( '-' converted to '+' ) \[\frac{a}{{x - \sqrt{y}}} \implies \frac{a}{{x - \sqrt{y}}} \times \frac{{x + \sqrt{y}}}{{x + \sqrt{y}}} \implies \frac{ax + a \sqrt{y}}{{x^2 - y}}\] Now here one more thing I want to tell you: You noticed that by using \({x + \sqrt{y}}\) as Rationalization Factor you will get \(x^2 - y\) : The y term is negative.. But in case of complex number when you rationalize in the same way, then you will get: \(x^2 + y\).. I show you: Take the same example but suppose you have complex number in denominator: \({x - i \cdot\sqrt{y}}\).. So: \[\frac{a}{{x - i \cdot \sqrt{y}}} \implies \frac{a}{{x - i \cdot\sqrt{y}}} \times \frac{{x + i \cdot \sqrt{y}}}{{x + i \cdot \sqrt{y}}} \implies \frac{ax + a \cdot i \cdot \sqrt{y}}{{x^2 - i^2 \cdot (\sqrt{y}})^2} \implies \frac{ax + a \cdot i \cdot \sqrt{y}}{{x^2 + y}}\] Since \(i^2 = -1 \) which makes it positive...

OpenStudy (lgbasallote):

isnt that called conjugation? is it still under rationalization? o.O

OpenStudy (anonymous):

Yes...

OpenStudy (lgbasallote):

hmm yup it is

OpenStudy (anonymous):

Rationalization basically means that to rationalize the denominator.. Don't you think we here making our denominator rational...??

OpenStudy (jiteshmeghwal9):

means making the denominator a whole number, isn't it?? @waterineyes

OpenStudy (anonymous):

Yes very right... Or it can be rational too..

OpenStudy (jiteshmeghwal9):

such as